OpenStudy (anonymous):

solve by substitution* -7x - y = -6 x - 2y = -12

OpenStudy (anonymous):

This one is easy. -7x-y=-6 x-2y=-12 We take value of x in terms of y from the second equation: x=2y-12 And substitute it in the first equation: -7(2y-12) - y =-6 -14y + 84 - y=-6 -15y +84=-6 84+6=15y 90=15y 90/15=y 6=y.......................First Answer. Substitute in equation 2: x-2(6)=-12 x-12=-12 So, x=0. There's our answer: x=0, y=6.